Score: 0 of 5 pts Instructor-created question The credit card with the Transaction Description ransaction Amount ransactions described on the right Previous balancel $6280 0 uses the average daily balance method to calculate interest. The annual interest rate is 18% of the average daily balance. Calculate parts a-c using the statement on March 12 C the right arch 1 Billing date arch 7 Charge: Restaurant $60.00 arch 21 Charge: Car Repairs $220 00 March 5 Payment 250.00 credit arge Groceries $100.00 March 31 End of billing period Payment Due Date April 9 a. Find the average daily balance for the billing period. Round to the nearest cent The average daily balance for the billing period is ) Round to the nearest cent as needed) b. Find the interest to be paid on April 1, the next billing date. Round to the nearest cent The interest to be paid on April 1 is S□ Use the answer from part a to find this answer Round to the nearest cent as needed ) C. Find the balance due on April 1 The balance due on April t is (Use the answer from part b to find this answer
Answer #1

a. Computation of average daily balance:
Average daily balance =Total of your balance from each day in the billing cycle / 31 days
Here, Total of your balance from each day in the billing cycle :

Number of days

Calculation

Amount

March 1 to March 5 = 5days

$6,280*5days

31,400

March 6 = 1day

(6,280-250)*1

6,030

March 7 to March 12= 6day

(6,030+60)*6

36,540

March 13 to March 21 = 9 days

(6,090+ 100)*9

55,710

March 22 to March 31 =10 days

(6,190+220)*10

64,100

TOTAL 31 DAYS

TOTAL

193,780

AVERAGE DAILY BALANCE

Total / 31 days

193,780/ 31 = 6,250.97

b) Interest to be paid on April 1:
Average daily balance * interest rate = $6,250.97*18% *1/12 = 93.76

c)Balance due on April 1 = 6,250.97 + 93.76 = $6,344.73